# First-order methods Almost Always Avoid Saddle points: The case of Vanishing Step-sizes

### v2

@@ 1 Introduction
Acknowledgement
We thank Nicolas Boumal and Andreea-Alexandra Musat for pointing out the mistake of in the proof of the main theorem, which requires the Hessian of $f$ to be invertible.
@@ 3 Intuition and Overview
For $\lambda>0$, the term (3) converges to zero if and only if $\sum_{t=0}^{\infty}\alpha_{t}=+\infty$ which is true if $\alpha_{t}$ is $\Omega\left(\frac{1}{t}\right)$. Moreover, for $\lambda=0$ it holds that the term (3) remains a constant (independently of the choice of stepsize $\alpha_{k}$) and for $\lambda<0$ it holds that the term (3) diverges for $\alpha_{t}$ to be $\Omega\left(\frac{1}{t}\right)$. Therefore, for $\alpha_{k}$ being $\Omega\left(\frac{1}{k}\right)$ we conclude that $\lim_{k\to\infty}x_{k}=\mathbf{0}$ whenever the initial point $\mathbf{x}_{0}$ satisfies $x_{0}^{i}=0$ ($i$-th coordinate of $\mathbf{x}_{0}$) for $\lambda_{i}\geq 0$.
Hence, if $\mathbf{x}_{0}\in E_{s}:=\textrm{span}(e_{1},\ldots,e_{j})$333$\{e_{1},...,e_{d}\}$ denote the classic orthogonal basis of $\mathbb{R}^{d}$., then $\mathbf{x}_{t}$ converges to the saddle point $x^{*}$ and if $\mathbf{x}_{0}$ has a component outside $E_{s}$ then gradient descent diverges. For the example above, the global stable set of $\mathbf{x}^{*}$ is the subspace $E_{s}$ which is of measure zero since $E_{s}$ is not full dimensional.
@@ 4 Stable Manifold Theorem for Time Non-homogeneous Dynamical Systems
Let $H$ be a $d\times d$ real diagonal matrix with at least one negative eigenvalue, i.e. $H=diag\{\lambda_{1},...,\lambda_{d}\}$ with $\lambda_{1}\geq\lambda_{2}\geq...\lambda_{s}>0\geq\lambda_{s+1}\geq...\geq\lambda_{d}$ and assume $\lambda_{d}<0$. Let $\eta(k,\mathbf{x})$ be a continuously differentiable function such that $\eta(k,\mathbf{0})=\mathbf{0}$ and for each $\epsilon>0$, there exists a neighborhood of $\mathbf{0}$ in which it holds
@@
    $D_{\mathbf{x}}g(k,\mathbf{x}^{*})=I-\alpha_{k}G,$ $G$ real diagonalizable with at least one negative eigenvalue;
@@
(since $\ \ \ (\text{since $\mathbf{x}_{k}=\varphi^{-1}(\mathbf{z}_{k})=Q^{-1}\mathbf{z}_{k}+\mathbf{x}^{*}$})$ (16)
@@
Fix a point $\mathbf{x}_{0}\in W^{s}(\mathcal{A}^{*})$. Since $\tilde{g}(k,0,\mathbf{x}_{0})\rightarrow\mathbf{x}^{*}\in\mathcal{A}^{*}$, there exists some nonnegative integer $T$ and all $t\geq T$, such that $\tilde{g}(t,0,\mathbf{x}_{0})\in\bigcup_{\mathbf{x}^{*}\in\mathcal{A}^{*}}U_{\mathbf{x}^{*}}=\bigcup_{i=1}^{\infty}U_{\mathbf{x}^{*}_{i}}.$ So $\tilde{g}(t,0,\mathbf{x}_{0})\in U_{\mathbf{x}^{*}_{i}}$ for some $\mathbf{x}^{*}_{i}\in\mathcal{A}^{*}$ and all $t\geq T$. This is equivalent to $\tilde{g}(T+k,T,\tilde{g}(T,0,\mathbf{x}_{0}))\in U_{\mathbf{x}^{*}_{i}}$ for all $k\geq 0$, and this implies that $\tilde{g}(T,0,\mathbf{x}_{0})\in\tilde{g}^{-1}(T+k,T,U_{\mathbf{x}^{*}_{i}})$ for all $k\geq 0$. And then we have $\tilde{g}(T,0,\mathbf{x}_{0})\in\bigcap_{k=0}^{\infty}\tilde{g}^{-1}(T+k,T,U_{\mathbf{x}^{*}_{i}}).$ Denote $S_{i,T}:=\bigcap_{k=0}^{\infty}\tilde{g}^{-1}(T+k,T,U_{\mathbf{x}^{*}_{i}})$ and the above relation is equivalent to $\mathbf{x}_{0}\in\tilde{g}^{-1}(T,0,S_{i,T})$. Take the union for all nonnegative integers $T$, we have $\mathbf{x}_{0}\in\bigcup_{T=0}^{\infty}\tilde{g}^{-1}(T,0,S_{i,T}).$ And union for all $i$ we obtain that $\mathbf{x}_{0}\in\bigcup_{i=1}^{\infty}\bigcup_{T=0}^{\infty}\tilde{g}^{-1}(T,0,S_{i,T})$ implying that $W^{s}(\mathcal{A}^{*})\subset\bigcup_{i=1}^{\infty}\bigcup_{T=0}^{\infty}\tilde{g}^{-1}(T,0,S_{i,T}).$ Since $S_{i,T}\subset W_{n}(\mathbf{x}^{*})$ from Corollary 4.2, and $W_{n}(\mathbf{x}^{*})$ has codimension at least 1. This implies that $S_{i,T}$ has measure 0. Since the image of set zero-measure set under diffeomorphism is of measure 0, and countable union of zero-measure sets is of measure 0, we obtain that $W^{s}(\mathcal{A}^{*})$ is of measure 0. ∎
@@ 5 Applications
In this section, we apply Theorem 4.1 (or its corollary 4.2) to the four of the most commonly used first-order methods and we prove that each one of them avoids strict saddle points even with vanishing stepsize $\alpha_{k}$ of order $\Omega\left(\frac{1}{k}\right)$.
@@ 5.1 Gradient Descent
[…]
@@ 5.2 Mirror Descent
  • $\nabla_{R}\Phi$ diverges on the relative boundary of $M$, that is $\lim_{x\rightarrow\partial M}||\nabla_{R}\Phi(x)||=\infty$.
@@ Appendix B Lyapunov-Perron Method
where $B_{k}$ and $C_{k}$ are diagonal as well and corresponding to stable and unstable subspaces of $I-\alpha_{k}H$ at 00. Using the same notation of denoting $A\left(m,n\right)$, we denote
@@
Next we prove that $T$ is a contraction map when choosing sequence in a small enough neighborhood around 00.
Take $\mathbb{B}(\delta)$ a small enough neighborhood around 00 such that the Lipschitz condition is satisfied. Let $u=\{u_{n}\}_{n\in\mathbb{N}}\subset\mathbb{B}(\delta)$ and $v=\{v_{n}\}_{n\in\mathbb{N}}\subset\mathbb{B}(\delta)$. Then we have
@@
[…]
@@
[…]
@@
[…]
@@
    If $S_{k}>\frac{1}{\lambda}$, or equivalently $01-\lambda S_{k}<$ 0, then $S_{k+1}-S_{k}<0$, and $S_{k}$ decreases until $S_{k_{1}}<\frac{1}{\lambda}$ for some $k_{1}\in\mathbb{N}$.
@@
    If $S_{k}<\frac{1}{\lambda}$, or equivalently $01-\lambda S_{k}>$ 0, then $S_{k+1}-S_{k}>0$, and $S_{k}$ increases until $S_{k}>\frac{1}{\lambda}$.
@@ 1. $\lim_{k\rightarrow\infty}(Tx)_{k+1}^{+}=0$:
implying that $S_{k}\rightarrow\infty$. So we conclude that $\lim_{k}S_{k}=0$, contradicting to the fact that $\lim_{k}S_{k}$ exists. So the $\lim_{k}S_{k}$ must be 00 if $S_{k}$ is of case 1.
If $S_{k}$ is of case 2, then immediately lim $\liminf S_{k}=0$. Suppose that lim $\limsup S_{k}>0$. Since $S_{k}$ decreases whenever $S_{k}>\frac{\left\|x_{k}\right\|}{\lambda}$ and $S_{k}$ increases whenever $S_{k}<\frac{\left\|x_{k}\right\|}{\lambda}$, we can find a subsequence $S_{k_{m}}$, with $S_{k_{m}-1}<S_{k_{m}}$, converging to lim $\limsup S_{k}$ as $m\rightarrow\infty$. But this is impossible since $S_{k_{m}-1}<\frac{\left\|x_{k}\right\|}{\lambda}$ and then $S_{k_{m}-1}\rightarrow 0$ as $m\rightarrow\infty$, which means $\lim_{m}\left|S_{k_{m}-1}-S_{k_{m}}\right|$ is positive, contradicting to the fact that $\lim_{k}\left|S_{k}-S_{k+1}\right|=0$. And thus, we have $\limsup_{k}S_{k}=0$, meaning that $\lim_{k}S_{k}=0$.
@@ 2. $\lim_{k\rightarrow\infty}(Tx)_{k+1}^{-}=0$:
$\left\|x_{k+1}^{-}\right\|=\left\|C_{k+1}x_{k}^{-}+\eta^{-}(k,x_{k})\right\|\geq\left|\left\|C_{k+1}x_{k}^{-}\right\|-\left\|\eta^{-}(k,x_{k})\right\|\right|.$
Since $\left\|\eta^{-}(k,x_{k})\right\|\rightarrow 0$ as $k\rightarrow\infty$ due to $\alpha_{k}$, and $\left\|C_{k+1}x_{k}^{-}\right\|\rightarrow\infty$ as $k\rightarrow\infty$ by assumption that $x_{k}^{-}$ is bounded away from 00. But this contradicts to the assumption $x_{k+1}^{-}$ is bounded in $U$. The proof completes. ∎

### v3

@@ 3 Intuition and Overview
For $\lambda>0$, the term (3) converges to zero if and only if $\sum_{t=0}^{\infty}\alpha_{t}=+\infty$ which is true if $\alpha_{t}$ is $\Omega\left(\frac{1}{t}\right)$. Moreover, for $\lambda=0$ it holds that the term (3) remains a constant (independently of the choice of stepsize $\alpha_{k}$) and for $\lambda<0$ it holds that the term (3) diverges for $\alpha_{t}$ to be $\Omega\left(\frac{1}{t}\right)$. Therefore, for $\alpha_{k}$ being $\Omega\left(\frac{1}{k}\right)$ we conclude that $\lim_{k\to\infty}x_{k}=\mathbf{0}$ whenever the initial point $\mathbf{x}_{0}$ satisfies $x_{0}^{i}=0$ (ii-th coordinate of $\mathbf{x}_{0}$) for $\lambda_{i}\geq 0$.
[…]
@@ 4 Stable Manifold Theorem for Time Non-homogeneous Dynamical Systems
Let $H$ be an invertible $d\times d$ real diagonal matrix with at least one negative eigenvalue, i.e. $H=diag\{\lambda_{1},...,\lambda_{d}\}$ with $\lambda_{1}\geq\lambda_{2}\geq...\lambda_{s}>0>\lambda_{s+1}\geq...\geq\lambda_{d}$ and assume $\lambda_{d}<0$. Let $\eta(k,\mathbf{x})$ be a continuously differentiable function such that $\eta(k,\mathbf{0})=\mathbf{0}$ and for each $\epsilon>0$, there exists a neighborhood of $\mathbf{0}$ in which it holds
@@
    $D_{\mathbf{x}}g(k,\mathbf{x}^{*})=I-\alpha_{k}G,$ $G$ invertible, real diagonalizable with at least one negative eigenvalue;
@@
$\ \ \ (\text{since $\mathbf{x}_{k}=\varphi^{-1}(\mathbf{z}_{k})=Q^{-1}\mathbf{z}_{k}+\mathbf{x}^{*}$})$ (16)
@@
Fix a point $\mathbf{x}_{0}\in W^{s}(\mathcal{A}^{*})$. Since $\tilde{g}(k,0,\mathbf{x}_{0})\rightarrow\mathbf{x}^{*}\in\mathcal{A}^{*}$, there exists some nonnegative integer $T$ and all $t\geq T$, such that $\tilde{g}(t,0,\mathbf{x}_{0})\in\bigcup_{\mathbf{x}^{*}\in\mathcal{A}^{*}}U_{\mathbf{x}^{*}}=\bigcup_{i=1}^{\infty}U_{\mathbf{x}^{*}_{i}}.$ So $\tilde{g}(t,0,\mathbf{x}_{0})\in U_{\mathbf{x}^{*}_{i}}$ for some $\mathbf{x}^{*}_{i}\in\mathcal{A}^{*}$ and all $t\geq T$. This is equivalent to $\tilde{g}(T+k,T,\tilde{g}(T,0,\mathbf{x}_{0}))\in U_{\mathbf{x}^{*}_{i}}$ for all $k\geq 0$, and this implies that $\tilde{g}(T,0,\mathbf{x}_{0})\in\tilde{g}^{-1}(T+k,T,U_{\mathbf{x}^{*}_{i}})$ for all $k\geq 0$. And then we have $\tilde{g}(T,0,\mathbf{x}_{0})\in\bigcap_{k=0}^{\infty}\tilde{g}^{-1}(T+k,T,U_{\mathbf{x}^{*}_{i}}).$ Denote $S_{i,T}:=\bigcap_{k=0}^{\infty}\tilde{g}^{-1}(T+k,T,U_{\mathbf{x}^{*}_{i}})$ and the above relation is equivalent to $\mathbf{x}_{0}\in\tilde{g}^{-1}(T,0,S_{i,T})$. Take the union for all nonnegative integers $T$, we have $\mathbf{x}_{0}\in\bigcup_{T=0}^{\infty}\tilde{g}^{-1}(T,0,S_{i,T}).$ And union for all $i$ we obtain that $\mathbf{x}_{0}\in\bigcup_{i=1}^{\infty}\bigcup_{T=0}^{\infty}\tilde{g}^{-1}(T,0,S_{i,T})$ implying that $W^{s}(\mathcal{A}^{*})\subset\bigcup_{i=1}^{\infty}\bigcup_{T=0}^{\infty}\tilde{g}^{-1}(T,0,S_{i,T}).$ Since $S_{i,T}\subset W_{n}(\mathbf{x}^{*})$ from Corollary 4.2, and $W_{n}(\mathbf{x}^{*})$ has codimension at least 1. This implies that $S_{i,T}$ has measure 0. Since the image of set zero-measure set under diffeomorphism is of measure 0, and countable union of zero-measure sets is of measure 0, we obtain $W^{s}(\mathcal{A}^{*})$ is of measure 0. ∎
@@ 5 Applications
In this section, we apply Theorem 4.1 (or its corollary 4.2) to the four of the most commonly used first-order methods and we prove that each one of them avoids strict saddle points even with vanishing stepsize $\alpha_{k}$ of order $\Omega\left(\frac{1}{k}\right)$. We also assume that the Hessian $\nabla^{2}f(\mathbf{x}^{*})$ is invertible.
@@ 5.1 Gradient Descent
$\theta(k,\mathbf{x})-\theta(k,\mathbf{y})=\int_{0}^{1}\frac{d}{dt}\theta(k,t\mathbf{x}+(1-t)\mathbf{y})dt=\int_{0}^{1}D_{\mathbf{z}}\theta(k,\mathbf{z})|_{\mathbf{z}=t\mathbf{x}+(1-t)\mathbf{y}}\cdot(\mathbf{x}-\mathbf{y})dt.$
@@ 5.2 Mirror Descent
[…]
@@ 6 Conclusion
Acknowledgement
We thank Nicolas Boumal and Andreea-Alexandra Musat for pointing out that the proof of the main theorem to work requires the extra assumption that the Hessian of $f$ is invertible.
@@ Appendix B Lyapunov-Perron Method
where $B_{k}$ and $C_{k}$ are diagonal as well and corresponding to stable and unstable subspaces of $I-\alpha_{k}H$ at 0. Using the same notation of denoting $A\left(m,n\right)$, we denote
@@
Next we prove that $T$ is a contraction map when choosing sequence in a small enough neighborhood around 0.
Take $\mathbb{B}(\delta)$ a small enough neighborhood around 0 such that the Lipschitz condition is satisfied. Let $u=\{u_{n}\}_{n\in\mathbb{N}}\subset\mathbb{B}(\delta)$ and $v=\{v_{n}\}_{n\in\mathbb{N}}\subset\mathbb{B}(\delta)$. Then we have
@@
Use spectrum norm $\left\|\cdot\right\|$ for matrices, we have
@@
$\int_{k}^{\infty}\exp\left(-t^{1-p}\right)dt$ $=\int_{k^{1-p}}^{\infty}\exp(-u)\frac{1}{1-p}u^{\frac{1}{1-p}-1}du$ (90)
$=\frac{1}{1-p}\int_{k^{1-p}}^{\infty}\exp(-u)u^{\frac{1}{1-p}-1}du$ (91)
@@
which implies that $\int_{k}^{\infty}\exp(-t^{1-p})dt$ converges, so does the series 89.
@@
    If $S_{k}>\frac{1}{\lambda}$, or equivalently $1-\lambda S_{k}<0$, then $S_{k+1}-S_{k}<0$, and $S_{k}$ decreases until $S_{k_{1}}<\frac{1}{\lambda}$ for some $k_{1}\in\mathbb{N}$.
@@
    If $S_{k}<\frac{1}{\lambda}$, or equivalently $1-\lambda S_{k}>0$, then $S_{k+1}-S_{k}>0$, and $S_{k}$ increases until $S_{k}>\frac{1}{\lambda}$.
@@ 1. $\lim_{k\rightarrow\infty}(Tx)_{k+1}^{+}=0$:
implying that $S_{k}\rightarrow\infty$. So we conclude that $\lim_{k}S_{k}=0$, contradicting to the fact that $\lim_{k}S_{k}$ exists. So the $\lim_{k}S_{k}$ must be 0 if $S_{k}$ is of case 1.
If $S_{k}$ is of case 2, then immediately $\liminf S_{k}=0$. Suppose that $\limsup S_{k}>0$. Since $S_{k}$ decreases whenever $S_{k}>\frac{\left\|x_{k}\right\|}{\lambda}$ and $S_{k}$ increases whenever $S_{k}<\frac{\left\|x_{k}\right\|}{\lambda}$, we can find a subsequence $S_{k_{m}}$, with $S_{k_{m}-1}<S_{k_{m}}$, converging to $\limsup S_{k}$ as $m\rightarrow\infty$. But this is impossible since $S_{k_{m}-1}<\frac{\left\|x_{k}\right\|}{\lambda}$ and then $S_{k_{m}-1}\rightarrow 0$ as $m\rightarrow\infty$, which means $\lim_{m}\left|S_{k_{m}-1}-S_{k_{m}}\right|$ is positive, contradicting to the fact that $\lim_{k}\left|S_{k}-S_{k+1}\right|=0$. And thus, we have $\limsup_{k}S_{k}=0$, meaning that $\lim_{k}S_{k}=0$.
@@ 2. $\lim_{k\rightarrow\infty}(Tx)_{k+1}^{-}=0$:
[…]
Since $\left\|\eta^{-}(k,x_{k})\right\|\rightarrow 0$ as $k\rightarrow\infty$ due to $\alpha_{k}$, and $\left\|C_{k+1}x_{k}^{-}\right\|\rightarrow\infty$ as $k\rightarrow\infty$ by assumption that $x_{k}^{-}$ is bounded away from 0. But this contradicts to the assumption $x_{k+1}^{-}$ is bounded in $U$. The proof completes. ∎